### v2

# Reinforcement learning control of quantum error correction

Abstract
The promise of fault-tolerant quantum computing is challenged by environmental drift that relentlessly degrades the quality of quantum operations. The contemporary solution, halting the entire quantum computation for recalibration, is unsustainable for the long runtimes of the future algorithms [reiher2017elucidating, gidney2025factor]. We address this challenge by unifying calibration with computation, granting the quantum error correction process [ryan2021realization, krinner2022realizing, sivak2023real, acharya2024quantum, bluvstein2024logical, bluvstein2025architectural, lacroix2025scaling] a dual role: its error detection events are not only used to correct the logical quantum state, but are also repurposed as a learning signal, teaching a reinforcement learning agent [silver2017mastering, mnih2015human, levine2016end, shalev2016safe, ouyang2022training] to continuously steer the physical control parameters and stabilize the quantum system during the computation. We experimentally demonstrate this framework on a superconducting processor, improving the logical error rate stability of the surface code 3.5-fold against injected drift and pushing the performance beyond what is achievable with state-of-the-art traditional calibration and human-expert tuning. Simulations of surface codes up to distance-15 confirm the scalability of our method, revealing an optimization speed that is independent of the system size. This work thus enables a new paradigm: a quantum computer that learns to self-improve directly from its errors and never stops computing.
Quantum computers are fundamentally analog machines, which makes them extremely fragile compared to digital classical computers [landauer1995quantum]. Quantum error correction (QEC) is the leading strategy to overcome this vulnerability by effectively digitizing the errors [shor1995scheme, nielsen2010quantum]: through repetitive error detection embedded in QEC protocols, spurious analog evolution branches into a sequence of “error” or “no error” events. These binary error detection signals can then be decoded to correct the logical quantum state, providing a mechanism for reaching low logical error rates required for practical applications.
However, the digitization of errors alone is insufficient, and the practical success of QEC is critically dependent on the analog control of the constituent qubits: QEC is effective only if the physical gate error rate is kept significantly below a certain threshold, around $10^{-3}-10^{-2}$ [stephens2014fault]. The process of precisely tuning the system’s control parameters to meet this prerequisite is known as calibration [khaneja2005optimal, kelly2018physical]. While traditional calibration techniques have been able to achieve the necessary precision [acharya2024quantum, bluvstein2025architectural], the analog nature of the control makes the system performance susceptible to drift. Consequently, the challenge expands from merely reaching deep below the QEC threshold to continuously maintaining this performance amidst the non-stationarity.
The solution to the drift problem used in prior experiments [ryan2021realization, krinner2022realizing, sivak2023real, acharya2024quantum, bluvstein2024logical, bluvstein2025architectural] is to halt the entire QEC process for intermittent system recalibration. However, this complete decoupling of computation and calibration represents a fundamental bottleneck for useful future algorithms that require continuous execution times on the order of days or weeks [reiher2017elucidating, gidney2025factor]. Although theoretical proposals based on logical swaps or code deformation [fang2024caliscalpel] have been put forward to reconcile this dichotomy, they entail a large overhead in the circuit runtime, footprint, and operational complexity. Yet, the very process of error detection embedded in the QEC algorithm offers a more direct solution, since errors from imperfect calibrations produce detectable syndromes just like all other errors. A pioneering attempt in Ref. [kelly2016scalable] to leverage this by engineering a direct feedback loop from error detection to physical control relied on a heuristic approach that is difficult to scale. In this work, we propose a different paradigm: repurposing error detection events as a learning signal for an artificial intelligence agent.
Our framework is based on reinforcement learning (RL), an approach that demonstrated remarkable success in automating solutions to complex control problems across diverse fields. It has powered systems that achieve superhuman performance in intricate games [silver2017mastering, mnih2015human], enabled significant advances in robotics [levine2016end, shalev2016safe], and has recently become integral to refining the behavior of large language models [ouyang2022training]. In the field of quantum control, RL was theoretically explored for various problems [fosel2018reinforcement, bukov2018reinforcement, fosel2021quantum, sivak2022model, metz2023self] and applied in experiments to improve the performance of isolated gates [ding2023high, baum2021experimental, reuer2023realizing] and bosonic codes [sivak2023real]. Building on these foundations, here we demonstrate the effectiveness of RL for the system-level challenge of calibrating and steering an error-corrected quantum processor.
We perform our experiments on the distance-5 surface and color codes implemented with superconducting circuits, focusing on a quantum memory algorithm, where the logical state is being preserved over time via repetitive application of QEC cycles. Our RL agent manages over a thousand control parameters which specify how an abstract QEC circuit is translated into analog waveforms that control the quantum system. The agent successfully steers the system against injected drift, improving the logical error rate (LER) stability $2.4$-fold, a figure that increases to $3.5$-fold with complementary decoder steering. Beyond counteracting drift, RL fine-tuning of an already well-calibrated processor achieves an additional $20\%$ suppression of the LER, pushing performance beyond the limits of traditional physics-based calibration and human-expert tuning to establish a new QEC record for both codes [acharya2024quantum, lacroix2025scaling]. The agent is able to reach high-performance even starting from randomized initial control parameters, suggesting the potential to augment or replace elements of the traditional calibration stack.
Having demonstrated RL control of today’s state-of-the-art QEC processor, we further showcase the future potential of this framework using numerical simulations. They establish the compatibility of RL control with uninterrupted quantum computation and its scalability to distance-15 surface code with tens of thousands of control parameters. Our approach is fundamentally general, requiring only error detection signals and tunable controls. Although demonstrated here using a planar QEC code implemented with superconducting circuits, it is directly applicable to any physical qubit modality and QEC architecture, including those with spatially nonlocal connectivity [Bravyi2024]. Our findings thus firmly establish reinforcement learning as a promising path towards automating the control of large-scale error-corrected quantum systems.
QEC optimization
using error detection events
The main guiding principle in the construction of fault-tolerant QEC circuits is to ensure that errors leave detectable signatures in the measurement outcomes of the code stabilizers. These signatures are most conveniently represented using the notion of detectors [gidney2021stim], defined as sets of measurements that have deterministic total parity in the absence of errors. A detection event, corresponding to flipped parity, signals occurrence of errors in a certain space-time region of the circuit, called the detecting region [mcewen2023relaxing], see Fig. 1(b). Detecting regions are usually sensitive to multiple kinds of errors, and a decoder is needed to infer the likely error patterns from any given pattern of detection events [higgott2023sparse, beni2025tesseract]. After the decoder’s correction of the logical state, the remaining error is quantified by the LER – the principal measure of quality of the QEC process.
To execute QEC, the parameters of the quantum gates are stored in the memory of a classical controller, see Fig. 1(a). These parameters are used by the compiler to convert an abstract QEC logical circuit to its physical implementation. In our learning-based approach, this process is modified: the controller deliberately applies small, simultaneous perturbations to all control parameters during the computation. These perturbations translate into subtle changes in the statistics of detection events. The goal of the learning algorithm is to disentangle cause and effect, correlating specific perturbations with changes in performance to continuously calibrate the system.
Figure 1: Overview of RL control. (a) Hierarchy of the feedback loops in control of an error-corrected quantum system. The low-level loop with analog control and readout signals (purple) occurs on a time scale of one QEC cycle; the logical algorithm’s digital feedback loop (green) occurs on a time scale of the decoding latency; the learning feedback loop (pink), presented in this work, is not synchronized with the lower levels, and occurs on a time scale determined by the relevant system drift. The indicated time scales are characteristic of superconducting circuits quantum computing platform [acharya2024quantum]. (b) A small space-time chunk of the QEC circuit for the repetition code, highlighting two overlapping detecting regions. (c) One iteration of the learning process. In each epoch, a batch of control policy candidates is sampled from the policy distribution. A certain number of QEC cycles is executed with each policy candidate (shades of red and blue). The acquired QEC data is used to compute rewards by estimating error detection rates for each detector. This information, indicating the relative performance of each policy candidate, is converted by the learning algorithm into a small gradient step of the policy distribution. Then, a new batch of policy candidates is sampled and the process repeats.
In principle, LER is the ideal global objective function for QEC calibration and optimization. However, using it directly faces fundamental obstacles to scalability, particularly in multi-qubit stabilizer codes such as the surface code, for the following reasons: (i) The LER is suppressed with the code distance $d$ as $\varepsilon_{\rm L}\propto\Lambda^{-d/2}$, where the error suppression factor $\Lambda>2$ has already been demonstrated in experiments with superconducting circuits [acharya2024quantum] and neutral atoms [bluvstein2025architectural]. Thus, accurately resolving the LER would require an exponentially large number of QEC cycles. (ii) The optimization involves a vast number of parameters – already more than a thousand in our $d=5$ experiment, and scaling as $O(d^{2})$ – which renders global optimization from a single scalar metric impractical. (iii) The LER is unsuitable for real-time calibration and steering during a quantum computation, as the logical state is generally unknown. While direct LER optimization has proven effective in specific contexts, such as for small bosonic codes [sivak2023real], these scaling challenges necessitate a different approach for large multi-qubit codes.
To address these limitations, we construct a surrogate objective function $C$, which is an efficiently-computable local proxy for LER that enables high-dimensional optimization, see Fig. 2(a). Our surrogate objective is defined as the average rate of detection events, $C=\hat{\mathbb{E}}[D]$, where $D$ denotes the detectors in the circuit, and the empirical expectation value is computed over a sufficient number of QEC cycles. This surrogate objective function is inspired by a simple scaling model of the surface code, in which $\Lambda=\varepsilon_{\rm th}/\varepsilon$ and $C\propto\varepsilon$, where $\varepsilon\ll\varepsilon_{\rm th}$ is the average physical error rate and $\varepsilon_{\rm th}$ is the QEC threshold. According to this model, the gradients of LER and of the surrogate objective are simply related as $\nabla\log\varepsilon_{\rm L}=(d+1)/2\cdot\nabla\log C$. To confirm that this relation is approximately satisfied in a real experimental setting, we sample Gaussian perturbations in the control parameter space and evaluate finite-difference partial derivatives, see Fig. 2(b). We observe good empirical agreement in the regime of small perturbations, which suffices for our RL approach that relies on Monte Carlo gradient estimation.
The surrogate objective immediately alleviates the limitation (iii). It also alleviates (i), since resolving $C$ to a constant relative accuracy requires a number of QEC cycles that scales as $O(\varepsilon^{-1})$ and is independent of the code distance. To alleviate (ii) and enable efficient high-dimensional optimization, we harness the sparse structure of the surrogate objective. We rely on the locality of detectors in the QEC circuit to narrow down the dependence of each term in $C$ to only a small subset of control parameters – those related to the gates within the respective detecting region. These relations can be conveniently represented using a factor graph, a bipartite graph in which the factor nodes correspond to detectors, and the variable nodes correspond to the system control parameters, see Fig. 2(c) and Supp. Mat. II for more information.
Figure 2: Optimization with surrogate objective. (a) While LER is the principal measure of quality of the QEC process, it is impractical to use during optimization, see main text. This motivates adoption of a surrogate objective $C$. (b) Finite-difference partial derivatives experimentally evaluated along random directions in the control parameter space confirm the linear relation between the gradient of the true and surrogate objectives, with proportionality coefficient $(d+1)/2$ (black line). (c) The surrogate objective allows us to effectively utilize the sparse dependence of error detection rates on the system control parameters, represented here as a factor graph. The detector nodes are connected to the learnable control parameters of the gates within their respective detecting regions. In our distance-5 surface code experiment, on average each detector node is connected to $302$ parameter nodes, and each parameter node to $18$ detector nodes.
Without harnessing the sparsity of the surrogate objective, our problem is intractable for standard off-the-shelf optimizers. Some algorithms are compatible with large-scale nonlinear sparse optimization problems [scipy2020] represented as factor graphs, but in practice additional considerations play an important role. Specifically, our objective function $C$ is highly stochastic, as it is obtained by averaging the binary detector values from a finite (and preferably small) number of QEC cycles. Furthermore, $C$ is non-stationary due to inherent system drift. Additionally, evaluating $C$ is expensive, since it requires updating the classical controller and acquiring data from the QEC detectors. Finally, in the future, $C$ must be optimized in real time during the quantum computation, favoring smooth and gradual improvements over disruptive large parameter steps that might lead to catastrophic consequences for the logical algorithm.
Given these properties, we adopted a multi-objective policy-gradient reinforcement learning approach (see Appendix G of Ref. [sivak2024optimization] for derivation), where the error detection rates of all constituent detectors serve as the multiple objectives optimized simultaneously. Our algorithm is built on parameter-exploring policy gradients [sehnke2010parameter], an approach where an entire control policy is sampled in one piece, which provides a convenient interface with the classical controller, see Fig. 1. For robustness in the stochastic and non-stationary setting, we integrate several key techniques: proximal policy optimization [schulman2017proximal] for stability, entropy regularization [haarnoja2018soft] to maintain exploration, and a replay buffer [mnih2015human] for improved sample efficiency. Crucially, the algorithm harnesses the sparse structure of the factor graph for efficient variance reduction of the Monte Carlo gradient estimator [mohamed2020monte] via gradient masking.
Our algorithm maintains a parameterized probability distribution over all system control parameters [sehnke2010parameter]. Each sample from this distribution, the so-called control policy, is a candidate solution to the optimization problem. As illustrated in Fig. 1(c), the learning proceeds iteratively, by repeating the same sequence of steps in every learning epoch. First, a batch of policy candidates is sampled and scored according to $C$: policies that lead to a lower rate of error detection events receive a higher reward. In practice, the detection rate (DR) is estimated by acquiring data from a finite number of QEC cycles executed with a given control policy, and then switching to another policy.
Next, the reward information, reflecting the relative performance of all policy candidates, is used by the learning algorithm to compute the gradient of the policy distribution, pushing it towards a more optimal region in the control parameter space. We use a factorized multivariate Gaussian distribution for simplicity. The mean $\mu(t)$ of this distribution represents the best guess of the optimal control policy during the learning epoch $t$. The diagonal covariance $\sigma(t)^{2}$ of the distribution controls the degree of exploration of the control space. It evolves during the learning process, typically shrinking over time in order to finely localize the solution [sivak2022model]. Eventually, the distribution converges to a local optimum of the optimization objective, in which detection event probabilities are minimized, indicating that occurrence of errors is suppressed to its minimum. In a non-stationary setting, the stochastic gradients steer the control parameters to follow the system drift. In this case, $\mu(t)$ learns to track the optimal policy over time, while $\sigma(t)^{2}$ maintains finite spread to never cease exploring.
RL fine-tuning of control policy
To demonstrate the practical utility of our RL framework, we first show that it can push system performance beyond the limits achievable by the state-of-the-art traditional calibration stack and human-expert tuning.
The central idea in the traditional approach to gate calibration is to orthogonalize the space of control parameters, using an accurate physical model of the system, and then construct dedicated “calibration experiments” that target a few parameters at a time. For example, a control parameter corresponding to the frequency of a microwave pulse for XY-rotation of a qubit is calibrated via spectroscopy, and then the amplitude of this pulse is calibrated via Rabi oscillations. Control parameters can also be jointly fine-tuned in calibration experiments that measure an aggregate performance metric of the gate, such as randomized or cross-entropy benchmarking [knill2008randomized, arute2019quantum]. Recently, this approach has culminated in a framework for automated calibration of large quantum systems based on the traversal of a directed acyclic graph [kelly2018physical], which encodes the order of calibrations of various control parameters, distilled from decades of research in the field of quantum control [koch2022quantum]. It has been successfully applied by the Google Quantum AI team in numerous experiments [arute2019quantum, acharya2024quantum, lacroix2025scaling], reaching the scale of $O(10^{2})$ qubits, as well as by other research teams [ding2023high, wu2021strong].
Using this traditional calibration approach and extensive human-expert tuning, we calibrate a distance-5 surface code and color code realized in a Willow quantum processor based on superconducting circuits [acharya2024quantum]. Subsequent RL fine-tuning of the system consistently yields about $20\%$ additional suppression of the LER after a few hundred training epochs, see Fig. 3(a). We achieve record distance-5 QEC performance across any physical qubit modality, for both the surface code and color code, see Fig. 3(b). The Tesseract most-likely-error decoder [beni2025tesseract] with a device-independent prior yields the LER of $\varepsilon_{L}=0.92\times 10^{-2}$ for the color code and $\varepsilon_{L}=1.92\times 10^{-3}$ for the surface code. Notably, under equivalent decoding the performance of our distance-5 surface code nearly matches that of the distance-7 code in Ref. [acharya2024quantum]. We expect that in the same system the gains from RL could be further enhanced by incorporating additional control parameters into its scope, e.g. the parameters of readout and reset gates. Beyond fine-tuning, in Supp. Mat. IV we show that RL successfully recovers the performance after we artificially randomize the control policy to fully scramble the logical observable, mimicking a coarse calibration scenario.
Generally, as gate error rates are pushed deeper below the QEC threshold, they will become limited by a growing number of low-probability error channels. These can be related to the unavoidable simplifying assumptions and approximations in the system models used by the standard calibration approach, or to yet unknown device physics. We anticipate that in the deep below-threshold regime, achieving and maintaining optimal system performance through metrology and targeted calibrations alone will become infeasible, and holistic model-free in-context fine-tuning of the QEC system will become requisite. Our results indicate that RL from error detection events could fill this critical role.
Figure 3: RL fine-tuning of QEC performance. (a) Systematic improvement of LER from RL fine-tuning applied after exhaustive conventional calibration process, with five independent runs for surface code and color code each (grey), mean performance (teal) and one-sigma deviation (shaded region). (b) Decay of the logical observable in a quantum memory experiment, averaged over $X$ and $Z$ bases, for surface code and color code. The reference curves (grey) use QEC syndrome data from Refs. [acharya2024quantum, lacroix2025scaling], here reprocessed with Tesseract decoder for consistent comparison, see main text. For Ref. [acharya2024quantum], we selected the best-performing among the distance-5 codes. Note, better results were achieved in Refs. [acharya2024quantum, lacroix2025scaling] using a more accurate neural network decoder, see Supp. Mat. VI.
RL steering of control policy
Next, we demonstrate that RL is able to not only optimize the system performance, but also steer the control policy in the presence of drift. To systematically study this effect, we inject artificial drift on several control parameters simultaneously, shown with red symbols in Fig. 4(a). We choose different temporal drift profiles (step-like, sinusoidal, and stroboscopic), different control parameters (CZ coupling strength, XY pulse amplitude, and frequency), and different locations of drifting gates on the qubit grid.
Following the already-described methodology, we start by calibrating the control policy at $t=0$. The performance of this policy (maroon) degrades over time due to the injected drift, since the fixed values of the control parameters become “outdated”, causing additional errors. These control errors lead to an elevated detection event signal, which appears exactly in those detectors where it is expected based on the constructed factor graph, see tiles highlighted with colored background in Fig. 4(a). In contrast to a fixed control policy, RL steering (blue) maintains a significantly suppressed rate of error detection events that consistently remains below the initial level (white line), except for brief moments when the drift is too fast. In Fig. 4(b), we show the evolution of the learned control parameters. The recovery from a step-like drift in the XY pulse amplitude (red circle) allows us to estimate the response time of the steering process of about $130$ epochs. This also sets the time scale for the policy lag in case of slow continuous drift, as in XY pulse frequency (red diamond).
To confirm that suppression of detection events is not due to hindered detection capability but is in fact due to suppressed errors, we evaluate the logical performance in Fig. 4(c). Compared to the fixed policy, we find on average a $24\%$ reduction of LER and a 2.42.4x improvement of its stability (quantified by the standard deviation of the LER distribution). These figures of merit further improve to $31\%$ and 3.53.5x respectively, by additionally steering the decoder parameters, as schematically illustrated in Fig. 1(a). Decoder steering is achieved within the same RL framework by reweighting the matching graph as described in Ref. [sivak2024optimization]. While steering of the classical controller is done via the surrogate objective $C$ that relies exclusively on the error detection probabilities, our decoder steering process relies on LER estimation, which is not straightforwardly scalable to the real-time setting. However, alternative approaches to decoder steering have been proposed [google2021exponential, wang2023dgr, hockings2025improving] that, in principle, do not suffer from this limitation.
We also analyzed the RL performance under natural system drift, which arises from sources ranging from material defects near the quantum system [klimov2018fluctuations] to temperature fluctuations in the classical control instruments and, unlike our prior demonstration, rarely has a simple time-dependence. Fourier analysis of multiple experimental RL runs from Fig. 3(a) reveals that the effect of steering can be understood as a filter that provides about $6\;\rm dB$ of suppression to low-frequency LER fluctuations originating from these natural sources, see Supp. Mat. III.
Figure 4: Demonstration of RL steering. (a) The data qubits (gold diamonds) and measure qubits (panels with data) are arranged in the layout of a distance-5 surface code. We inject artificial drift on the gates indicated with red shapes (circle, diamond, bars) and observe elevated error detection signals where expected (colored background). The detection rate associated with each measure qubit is normalized for visualization to remove the effect of the natural system drift. Performance of the fixed control policy (maroon) degrades over time due to the injected drift, while RL steering (blue) stabilizes and maintains the DR below its initial level (white lines). (b) Time-dependence of the injected drift in system control parameters (dashed) and RL steering (solid). Exponential fit of RL response to step-like drift in XY pulse amplitude yields the characteristic learning time of $130$ epochs. (c) Periodic evaluation of the logical performance indicates that RL steering of the system significantly suppresses and stabilizes the LER, see main text. Additionally incorporating the decoder steering (black) further improves these results.
We have thus far demonstrated that the mean $\mu(t)$ of the Gaussian policy distribution learns to track the optimal policy over time in the presence of drift. As a result, the learned policy $\mu(t)$ has superior performance compared to the fixed policy, with lower DR and LER. However, since the learning process requires exploration of the parameter space, the algorithm inevitably samples policy candidates whose performance is worse than that of the $\mu(t)$ policy. This “exploration noise” is irrelevant in our experimental setting where RL steering relies on repetitive executions of a short logical algorithm and the quantum state is independently re-prepared in every shot. However, in the future this steering must be done in real time during the single-shot execution of a long logical algorithm. In that case, the exploration noise, although necessary for learning, is going to be detrimental to the logical algorithm’s performance. Generally, balancing the exploration of parameter space and exploitation of the learned policy $\mu(t)$ is a central challenge in many applications of RL [sutton2018reinforcement]. In our case, the favorable resolution of the exploration-exploitation tradeoff would mean that aggregate performance of all sampled policy candidates, most of which are worse than $\mu(t)$, is still better than the performance without steering.
To study this tradeoff, we conducted numerical simulations of real-time steering of the distance-3 surface code subject to sinusoidal parameter drift at different frequencies, see Fig. 5(a) and Supp. Mat. V A. We count the total number of error detection events generated in a $1.8\times 10^{9}$-cycle window of QEC, and normalize it so that level 1 corresponds to performance of the optimal policy (known in the simulation), and level 0 corresponds to the performance of a fixed policy (optimal at $t=0$). We control the exploration-exploitation tradeoff by changing the amount of entropy regularization [haarnoja2018soft] in our RL algorithm. Our findings indicate that there is a critical drift frequency, about $1/150$ epochs, below which the system becomes real-time steerable: the performance with exploration noise is better than the performance of a fixed policy. This critical frequency is consistent with the response time of the learning algorithm observed in experiment, see Fig. 4(b). When drift is too fast, the real-time steering is not able to keep up – such drift must be mitigated at the hardware level. In contrast, at low drift frequencies, the exploration and exploitation can be successfully balanced to closely approach the performance of the optimal policy at all times.
Thus, our simulations establish that RL is able to effectively utilize the information concealed in the error detection events to calibrate the system while continuing the logical computation. This ability offers a significant advantage over approaches based on the synthesis of traditional calibration and code deformation [fang2024caliscalpel], since RL steering does not introduce any resource overhead.
Scaling and outlook
Figure 5: Real-time steering and scaling simulations. (a) Normalized improvement (color) of detection rate in the real-time steering simulation of the distance-3 surface code subject to sinusoidal drift at different frequencies. Level 11 indicates the performance of optimal policy. Isoline at level 0 demarcates the boundary beyond which real-time steering results in better performance than a fixed policy, approaching the performance of optimal policy in the regime of slow drift. (b) Simulation of scalability of RL control of large surface codes. The algorithm learns the parameters of single-qubit and CZ gates, with $30$ control parameters per gate, amounting to almost $40,000$ control parameters in total for the distance-15 code. During the learning process, the logical error rate reduces over time (color) until it reaches the floor (red bars) set by the irreducible physical error rates and characterized by the error suppression factor $\Lambda^{*}$. (c) Point-estimates of $\Lambda$ at every code distance and learning epoch from (b) confirm that the speed, $\partial_{t}\Lambda/\Lambda^{*}\times 10^{2}$, at which the error suppression factor approaches the local optimum, is proportional to the distance from the optimum, $1-\Lambda/\Lambda^{*}$. The convergence rate $\gamma$ is independent of the system size but depends on the number of control parameters per gate, with three beams corresponding to 1, 10, and 30 parameters, and the linear fits (red) indicating the convergence rates.
Finally, we demonstrate the scalability of RL control with simulations extending to a distance-15 surface code with approximately $40,000$ control parameters, see Fig. 5(b). We extract point-estimates of the “instantaneous” error suppression factor $\Lambda$ at each code distance and learning epoch, verifying in Fig. 5(c) that its convergence rate $\partial_{t}\Lambda/\Lambda^{*}$ is proportional to the distance from the optimal $\Lambda^{*}$, set in our simulation by the irreducible control-independent physical error rates. This empirical result leads to the exponential convergence law, $1-\Lambda/\Lambda^{*}\propto e^{-\gamma\,t}$, characteristic of gradient-descent based algorithms in the vicinity of local optimum [boyd2004convex], with the convergence rate $\gamma$ that depends on properties such as the number of gates per detecting region and the number of control parameters per gate. Crucially, this rate is independent of the system size, a direct consequence of our algorithm’s ability to effectively utilize the sparsity of the factor graph representation of the control problem.
Multiple possible extensions of our RL control framework leave ample room to further enhance its capabilities. For example, instead of relying on a simple Gaussian policy distribution, it could be advantageous to employ deep neural networks [mnih2015human] and condition the policy distribution on the observations from the environment, e.g. on some suitable statistics of the error detection events. Learning the system model in addition to the policy [hafner2019dream] can lead to better sample efficiency and further improve the quality of solutions by discovering unexpected correlations in the data. It is conceivable that in the future, with sufficient enhancement of our learning framework, a quantum processor could be calibrated for QEC ab initio fully by RL, with no reliance on the traditional calibration paradigm or human experts.
In conclusion, by empowering the quantum computer to learn directly from its errors, we unlock a scalable pathway to optimize performance in real time, replacing disruptive calibration routines with uninterrupted computation. Our work suggests that the path to fault tolerance will be built not only on better hardware, but on more intelligent control.
Contributions
V.S. conceived and led the project, with significant contribution from A.M; V.S. developed the RL control framework; A.M. calibrated and expert-tuned the processor; V.S. and A.M. designed and conducted the experiments and wrote the manuscript with input from all the authors; M.B. and M.N. contributed software improvements; A.E. performed device simulations; P.K. supervised the project. All authors contributed to building and maintaining the hardware, software, cryogenics, and electronics infrastructure.
Corresponding authors
Correspondence and requests for materials should be addressed to Volodymyr Sivak (vladsivak@google.com) and Alexis Morvan (amorvan@google.com).
Competing Interests
The authors declare no competing interests.
Code and data availability
The data supporting our findings are available on Zenodo at https://doi.org/10.5281/zenodo.17566522.
Google Quantum AI and Collaborators
Volodymyr Sivak1, Alexis Morvan1, Michael Broughton1, Matthew Neeley1, Alec Eickbusch1, Dmitry Abanin1, Amira Abbas1, Rajeev Acharya1, Laleh Aghababaie Beni1, Georg Aigeldinger1, Ross Alcaraz1, Sayra Alcaraz1, Trond I. Andersen1, Markus Ansmann1, Frank Arute1, Kunal Arya1, Walt Askew1, Nikita Astrakhantsev1, Juan Atalaya1, Brian Ballard1, Joseph C. Bardin1, 2, Hector Bates1, Andreas Bengtsson1, Majid Bigdeli Karimi1, Alexander Bilmes1, Simon Bilodeau1, Felix Borjans1, Alexandre Bourassa1, Jenna Bovaird1, Dylan Bowers1, Leon Brill1, Peter Brooks1, David A. Browne1, Brett Buchea1, Bob B. Buckley1, Tim Burger1, Brian Burkett1, Nicholas Bushnell1, Jamal Busnaina1, Anthony Cabrera1, Juan Campero1, Hung-Shen Chang1, Silas Chen1, Ben Chiaro1, Liang-Ying Chih1, Agnetta Y. Cleland1, Bryan Cochrane1, Matt Cockrell1, Josh Cogan1, Roberto Collins1, Paul Conner1, Harold Cook1, Rodrigo G. Cortiñas1, William Courtney1, Alexander L. Crook1, Ben Curtin1, Martin Damyanov1, Sayan Das1, Dripto M. Debroy1, Sean Demura1, Paul Donohoe1, Ilya Drozdov1, Andrew Dunsworth1, Valerie Ehimhen1, Aviv Moshe Elbag1, Lior Ella1, Mahmoud Elzouka1, David Enriquez1, Catherine Erickson1, Vinicius S. Ferreira1, Marcos Flores1, Leslie Flores Burgos1, Ebrahim Forati1, Jeremiah Ford1, Austin G. Fowler1, Brooks Foxen1, Masaya Fukami1, Alan Wing Lun Fung1, Lenny Fuste1, Suhas Ganjam1, Gonzalo Garcia1, Christopher Garrick1, Robert Gasca1, Helge Gehring1, Robert Geiger1, Élie Genois1, William Giang1, Dar Gilboa1, James E. Goeders1, Edward C. Gonzales1, Raja Gosula1, Stijn J. de Graaf1, Alejandro Grajales Dau1, Dietrich Graumann1, Joel Grebel1, Alex Greene1, Jonathan A. Gross1, Jose Guerrero1, Loïck Le Guevel1, Tan Ha1, Steve Habegger1, Tanner Hadick1, Ali Hadjikhani1, Michael C. Hamilton1, Matthew P. Harrigan1, Sean D. Harrington1, Jeanne Hartshorn1, Stephen Heslin1, Paula Heu1, Oscar Higgott1, Reno Hiltermann1, Hsin-Yuan Huang1, Mike Hucka1, Christopher Hudspeth1, Ashley Huff1, William J. Huggins1, Evan Jeffrey1, Shaun Jevons1, Zhang Jiang1, Xiaoxuan Jin1, Chaitali Joshi1, Pavol Juhas1, Andreas Kabel1, Dvir Kafri1, Hui Kang1, Kiseo Kang1, Amir H. Karamlou1, Ryan Kaufman1, Kostyantyn Kechedzhi1, Tanuj Khattar1, Mostafa Khezri1, Seon Kim1, Can M. Knaut1, Bryce Kobrin1, Fedor Kostritsa1, John Mark Kreikebaum1, Ryuho Kudo1, Ben Kueffler1, Arun Kumar1, Vladislav D. Kurilovich1, Vitali Kutsko1, Nathan Lacroix1, David Landhuis1, Tiano Lange-Dei1, Brandon W. Langley1, Pavel Laptev1, Kim-Ming Lau1, Justin Ledford1, Joy Lee1, Kenny Lee1, Brian J. Lester1, Wendy Leung1, Lily Li1, Wing Yan Li1, Ming Li1, Alexander T. Lill1, William P. Livingston1, Matthew T. Lloyd1, Aditya Locharla1, Laura De Lorenzo1, Daniel Lundahl1, Aaron Lunt1, Sid Madhuk1, Aniket Maiti1, Ashley Maloney1, Salvatore Mandrà1, Leigh S. Martin1, Orion Martin1, Eric Mascot1, Paul Masih Das1, Dmitri Maslov1, Melvin Mathews1, Cameron Maxfield1, Jarrod R. McClean1, Matt McEwen1, Seneca Meeks1, Kevin C. Miao1, Zlatko K. Minev1, Reza Molavi1, Sebastian Molina1, Shirin Montazeri1, Charles Neill1, Michael Newman1, Anthony Nguyen1, Murray Nguyen1, Chia-Hung Ni1, Murphy Yuezhen Niu1, Logan Oas1, Raymond Orosco1, Kristoffer Ottosson1, Alice Pagano1, Agustin Di Paolo1, Sherman Peek1, David Peterson1, Alex Pizzuto1, Elias Portoles1, Rebecca Potter1, Orion Pritchard1, Michael Qian1, Chris Quintana1, Arpit Ranadive1, Matthew J. Reagor1, Rachel Resnick1, David M. Rhodes1, Daniel Riley1, Gabrielle Roberts1, Roberto Rodriguez1, Emma Ropes1, Lucia B. De Rose1, Eliott Rosenberg1, Emma Rosenfeld1, Dario Rosenstock1, Elizabeth Rossi1, Pedram Roushan1, David A. Rower1, Robert Salazar1, Kannan Sankaragomathi1, Murat Can Sarihan1, Kevin J. Satzinger1, Max Schaefer1, 3, Sebastian Schroeder1, Henry F. Schurkus1, Aria Shahingohar1, Michael J. Shearn1, Aaron Shorter1, Noah Shutty1, Vladimir Shvarts1, Spencer Small1, W. Clarke Smith1, David A. Sobel1, Barrett Spells1, Sofia Springer1, George Sterling1, Jordan Suchard1, Aaron Szasz1, Alexander Sztein1, Madeline Taylor1, Jothi Priyanka Thiruraman1, Douglas Thor1, Dogan Timucin1, Eifu Tomita1, Alfredo Torres1, M. Mert Torunbalci1, Hao Tran1, Abeer Vaishnav1, Justin Vargas1, Sergey Vdovichev1, Guifre Vidal1, Catherine Vollgraff Heidweiller1, Meghan Voorhees1, Steven Waltman1, Jonathan Waltz1, Shannon X. Wang1, Brayden Ware1, James D. Watson1, Yonghua Wei1, Travis Weidel1, Theodore White1, Kristi Wong1, Bryan W. K. Woo1, Christopher J. Wood1, Maddy Woodson1, Cheng Xing1, Z. Jamie Yao1, Ping Yeh1, Bicheng Ying1, Juhwan Yoo1, Noureldin Yosri1, Elliot Young1, Grayson Young1, Adam Zalcman1, Ran Zhang1, Yaxing Zhang1, Ningfeng Zhu1, Nicholas Zobrist1, Zhenjie Zou1, Ryan Babbush1, Dave Bacon1, Sergio Boixo1, Yu Chen1, Zijun Chen1, Michel Devoret1, 3, Monica Hansen1, Jeremy Hilton1, Cody Jones1, Julian Kelly1, Alexander N. Korotkov1, Erik Lucero1, Anthony Megrant1, Hartmut Neven1, William D. Oliver1, Ganesh Ramachandran1, Vadim Smelyanskiy1, Paul V. Klimov1
$^{\textrm{\scriptsize 1}}$ Google Quantum AI
$^{\textrm{\scriptsize 2}}$ Department of Electrical and Computer Engineering, University of Massachusetts, Amherst, MA
$^{\textrm{\scriptsize 3}}$ Department of Physics, University of California, Santa Barbara, CA